A man of mass 50kg States moving on the earth and acquires a speed of 1.8 m/sWith what speed does the earth recoil ? Mass of the earth is 6×10²⁴ kg

Grade:12th pass

1 Answers

RITWIK RAJ JHA
7 years ago
Momentum of man=50×1.8=90Now by conservation of momentum theorem =90=6×10^24×x (where x is velocity of recoilation of earth).so